OpenStudy (anonymous):

x2 – 8x + 15 a2 – a – 20 a2 – 5a – 20 a2 + 12ab + 27b2 2a2 + 30a + 100

OpenStudy (nottim):

wat we do

OpenStudy (akshay_budhkar):

lol

hero (hero):

funny

OpenStudy (akshay_budhkar):

should we add it for you?

OpenStudy (nottim):

Woo. nice work everyone.

OpenStudy (akshay_budhkar):

the most probable conclusion is that the user wants us to factorise it

OpenStudy (anonymous):

si

OpenStudy (anonymous):

<3:)

OpenStudy (akshay_budhkar):

tht is yes in french ^

OpenStudy (nottim):

woo. victory.

OpenStudy (akshay_budhkar):

ok i will do the first one for you x^2-3x-5x+15 =x(x-3)-5(x-3) =(x-3)(x-5)

OpenStudy (anonymous):

ty sir

OpenStudy (akshay_budhkar):

will you do the rest?

OpenStudy (anonymous):

UGHHHH fine, only for you

OpenStudy (anonymous):

can someone help me please after you are done here? http://openstudy.com/#/updates/4eeea895e4b0367162f5f40f

OpenStudy (akshay_budhkar):

lol if u want me to help i can, but its better you try it on your own right?

OpenStudy (anonymous):

i guess:(

OpenStudy (anonymous):

ill post again if i dont understand one

hero (hero):

Anybody who says uhhhmath is probably looking for answers anyway

OpenStudy (anonymous):

you very clever

OpenStudy (nottim):

i am

OpenStudy (akshay_budhkar):

lol

OpenStudy (anonymous):

so noone knows how to do my probelm?

OpenStudy (anonymous):

problem*

OpenStudy (anonymous):

=\

OpenStudy (akshay_budhkar):

i have probably never heard of that number,